### v1

@@ 10.3 Using Value Estimation to Inform Behaviour
SVERL’s value estimation explanations reveal how feature values influence the agent’s return estimate. As shown in the examples in Section 7, feature values in a given state that also appear in high-return states tend to receive positive attributions. In contrast, feature values associated with lower-return states tend to receive negative attributions. This suggests that these explanations are useful for highlighting which aspects of the environment correlate with successful estimated outcomes. One intriguing possibility is to use this information to inform behaviour—for example, encouraging actions that lead to positively attributed features, or identifying negatively attributed features that should be changed to improve expected return. The additional example in Section B.3 illustrates this idea concretely: in a dice-rolling domain, each die is treated as a feature and receives a negative attribution if and only if the optimal policy re-rolls the corresponding die to change its value. How to effectively incorporate these signals into decision-making or policy optimisation is a compelling question, and one we believe represents an important direction for future work.
10.4 Connecting SVERL to Broader Work on Shapley Values
@@
SVERL provides a principled framework for explaining reinforcement learning agents using Shapley values. It introduces new cooperative games tailored to reinforcement learning by defining policy, performance, and value estimation characteristic functions. Because the framework inherits the structure and axiomatic foundations of classic Shapley value theory, it can naturally accommodate alternative definitions, theoretical developments, and methodological advances from the broader literature.
@@
To illustrate the kind of opportunities this generality creates, we highlight two possible extensions. First, Frye et al. (2020b) propose computing Shapley values under a causal model, modifying the Shapley value sum to account for known dependencies between features. Extending SVERL in this way could yield explanations that more accurately reflect the causal structure of the environment or agent behaviour. Another possibility is to explore alternative solution concepts from cooperative game theory. For instance, Banzhaf values (Banzhaf III, 1964) define a different notion of attribution based on a distinct set of axioms and may offer complementary insights when applied within the SVERL framework. These examples demonstrate only two possible ways broader theoretical developments in Shapley value research can be incorporated into reinforcement learning using SVERL. Exploring these connections provides a promising avenue for future work.
@@ 11 Conclusion
Reinforcement learning has achieved impressive results across many domains, often surpassing human ability. Yet the how and why of agents interacting with their environments remains difficult to interpret, which hinders trust and limits adoption. In this work, we addressed this challenge by introducing Shapley Values for Explaining Reinforcement Learning (SVERL), a theoretical framework for explaining agent-environment interactions.
Motivated by an illustrative example, we identified three core elements of reinforcement learning to explain: behaviour, performance, and value estimation. Together, these offer a general understanding of reinforcement learning systems. We proposed a unified framework for explaining these elements through the influence of features at specific states. By treating features as players that cooperate to produce each element, we defined a family of theoretically principled explanations grounded in the well-established foundations of Shapley values in cooperative game theory.
We examined how these explanations relate to one another and how the framework as a whole can provide a more complete view of agent–environment interactions. We also discussed important considerations for interpreting SVERL in practice, including potential challenges and limitations. Finally, we considered possible extensions of the framework, such as explaining behaviour more broadly and approximating SVERL in large-scale settings.
By providing a theoretical foundation for explaining reinforcement learning, we aim to support the development of principled and interpretable reinforcement learning methods. We hope that this work lays the groundwork for future research towards more transparent and trustworthy systems.
@@
This work was supported by the UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI (ART-AI) [EP/S023437/1], the EPSRC Centre for Doctoral Training in Digital Entertainment (CDE) [EP/L016540/1] and the University of Bath. This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath.
@@ Appendix A Proofs of Propositions and Results
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the performance game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C}):2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the performance game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
@@
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the value estimation game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the estimated return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\hat{v}^{\pi}_{s}:2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the value estimation game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
@@
When comparing return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
@@ Appendix B Additional Examples
Additional examples are presented. We use three intuitive examples to illustrate how our framework attributes behaviour, performance and value estimates to features. We then conclude with an additional example explaining behaviour, performance, and value estimation.
B.1 Additional Example on Explaining Behaviour
State ($s$) Action ($a$) $p^{\pi}(s)$ $\tilde{\pi}_{s}^{a}(\text{Both})$ $\tilde{\pi}_{s}^{a}(\text{I})$ $\tilde{\pi}_{s}^{a}(\text{C})$ $\tilde{\pi}_{s}^{a}(\emptyset)$ $\phi_{\text{I}}$ $\phi_{\text{C}}$ $(1,\text{Red})$ N $0.25$ 00 00 00 $0.25$ $-0.125$ $-0.125$ E $1$ $1$ $1$ $0.25$ $0.375$ $0.375$ S 00 00 00 $0.25$ $-0.125$ $-0.125$ W 00 00 00 $0.25$ $-0.125$ $-0.125$ $(3,\text{Green})$ N $0.25$ $1$ $1$ $0.5$ $0.25$ $0.625$ $0.125$ E 00 00 00 $0.25$ $-0.125$ $-0.125$ S 00 00 00 $0.25$ $-0.125$ $-0.125$ W 00 00 $0.5$ $0.25$ $-0.375$ $0.125$
@@ B.1 Additional Example on Explaining Behaviour
We first consider explaining agent behaviour. Figure 6 shows a $2\times 2$ gridworld in which each state is defined by two features: its index and its colour. For example, the top-left state is indexed as 1, coloured red, and denoted $(1,\text{Red})$. The figure also shows a policy $\pi$ that moves clockwise through the grid: East in state $(1,\text{Red})$, South in state $(2,\text{Blue})$, West in state $(4,\text{Green})$, and North in state $(3,\text{Green})$. We analyse how the agent’s action probabilities in states $(1,\text{Red})$ and $(3,\text{Green})$ change as different features become known. These probabilities define SVERL’s characteristic values for explaining behaviour, from which the Shapley values are computed. Both are summarised in the table in Figure 6.
Consider first state $(1,\text{Red})$. When neither feature is known, the agent is equally likely to be in any state (by the steady-state distribution), and thus acts uniformly over the four actions. If the index becomes known to be $1$, or the colour becomes known to be Red, the agent must be in state $(1,\text{Red})$ and selects East with probability 1. Thus, both features fully resolve the agent’s uncertainty and have the same effect on the action probabilities. This is reflected in the Shapley values in Figure 6, which assign equal contributions to the index and colour for increasing the probability of the correct action and decreasing the probabilities of the incorrect ones.
Now consider state $(3,\text{Green})$. If the index becomes known to be $3$, the agent must be in state $(3,\text{Green})$ and selects North with probability 1. If instead the colour becomes known to be Green, the agent could be in either state $(3,\text{Green})$ or $(4,\text{Green})$, and hence selects North and West with equal probability. Thus, both features increase the probability of the correct action (North), the index increases this probability more, and the colour increases the probability of an incorrect action (West). This distinction is reflected in the Shapley values in Figure 6.
B.2 Additional Example on Explaining Performance
@@ 2 N 1 1 $2/3$ $6/7$
1 $1/7$ $6.00$ $6.00$ $4.00$ $0.00$ $4.00$ $2.00$
2 $2/7$ $7.00$ $7.00$ $6.50$ $6.83$ $0.33$ $-0.17$
Figure 8: A five-state gridworld with two goal states (G). The agent starts in state 1 or 2 with equal probability and follows the optimal deterministic policy (shown with arrows). Each state is defined by two features: its $x$- and $y$-coordinates. The top-right table shows the action probabilities under different feature subsets for states 1 and 2. The bottom table shows the characteristic values and Shapley attributions used to explain performance.
@@
We now consider explaining performance. Figure 8 shows a five-state gridworld where the agent begins uniformly at random in state 1 or 2, receiving a reward of $-1$ per action and an additional $+10$ upon reaching a goal state. The figure also shows the optimal deterministic policy $\pi$: East from state 1 and North from all others; invalid actions incur reward but leave the agent’s position unchanged. Each state is described by two features: its $x$- and $y$-coordinates. We analyse how the expected return from each state changes as these coordinates become known, owing to the agent’s behaviour in that state shifting accordingly. These returns define SVERL’s characteristic values for explaining performance, while the resulting Shapley values quantify the influence of each feature. Both are shown in the bottom table of Figure 8; the top-right table shows the corresponding action probabilities under different subsets of known features.
@@
Consider state 1. When neither feature is known, the agent has probability $1/7$ of being in state 1 (with optimal action East) and probability $2/7$ of being in states 2, 3, and 4 (all with optimal action North). Therefore, the agent selects North with probability $6/7$ and East with $1/7$, yielding an expected return of 00 from state 1. If the $x$-coordinate becomes known, the agent must be in state 1 and selects East, yielding a return of $6$. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability $2/3$ and East with $1/3$, giving an expected return of $4$. This difference is reflected in the Shapley values in Figure 8, which assign twice as much influence on expected return to the $x$-coordinate as the $y$-coordinate.
@@
Now consider state 2. As in state 1, when no features are known, the agent follows the steady-state distribution, selecting North with probability $6/7$ and East with probability $1/7$, yielding an expected return of $6.83$ from state 2. If the $x$-coordinate becomes known, the agent must be in states 2, 3, or 4 (all with optimal action North). Therefore, the agent deterministically selects North, increasing the return from state 2 to $7$. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability $2/3$ and East with $1/3$. This gives a lower expected return of $6.5$ because the likelihood of the optimal action (North) has decreased from $6/7$ to $2/3$. The Shapley values in Figure 8 reflect this: the $x$-coordinate contributes positively to performance, while the $y$-coordinate has a negative influence.
B.3 Additional Example on Explaining Value Estimation
Table 7: Characteristic values and Shapley values for value estimates in two states of a dice-rolling game. Each state is defined by two features: the values of two six-sided dice, denoted as $(d_{1},d_{2})$. The agent may re-roll either die, both dice, or neither. After each action, the episode ends with probability $0.5$, yielding a reward of $1$ if the sum of the dice exceeds $9$, and 00 otherwise. For the two states shown, the optimal policy re-rolls the $3$ in $(3,6)$ and re-rolls both dice in $(1,1)$.
@@ B.3 Additional Example on Explaining Value Estimation
$(3,6)$ $0.024$ $0.67$ $0.45$ $0.90$ $0.66$ $-0.22$ $0.23$
$(1,1)$ $0.018$ $0.36$ $0.45$ $0.45$ $0.66$ $-0.15$ $-0.15$
@@
Next, we consider explaining value estimation in a dice-rolling game. Each episode begins with a random roll of two six-sided dice. The resulting state is denoted $(d_{1},d_{2})$, where $d_{1}$ and $d_{2}$ are the first and second die values, respectively. The agent may re-roll either die, both dice, or neither. After each action, the episode ends with probability $0.5$, yielding a reward of $1$ if the sum of the dice is $10$ or more, and 00 otherwise. The optimal policy keeps $5$s and $6$s, and re-rolls $1$s, $2$s, $3$s, and $4$s. We analyse how the agent’s estimated return in each state changes as one or both dice become known. These returns define the characteristic values in SVERL’s value estimation game. The resulting Shapley values quantify each die’s contribution, as shown in Table 7.
Consider state $(3,6)$. When neither die is known, the value estimate is the average return over the steady-state distribution: $0.66$. If $d_{2}$ becomes known to be $6$, the estimate increases to $0.90$ because states with a $6$ have higher average returns—the agent keeps the $6$ and only possibly re-rolls the other die. If instead $d_{1}$ becomes known to be $3$, the estimate drops to $0.45$ because states with a $3$, which are re-rolled by the optimal policy, have lower average returns. The Shapley values reflect this difference: the $6$ contributes positively to the value estimate, while the $3$ contributes negatively.
Now consider state $(1,1)$. As above, when neither die is known, the estimate is $0.66$. If either die becomes known to be $1$, the estimate decreases to $0.45$ because states with a $1$ have a lower average return—the $1$ must be re-rolled. When both dice are known, the agent must be in the lowest-return state, and the estimate falls further to $0.36$. Thus, each die lowers the value estimate by the same amount, as reflected in their equal negative Shapley values in Table 7.
Figure 9: Shapley values for the value estimates of every state in the dice-rolling game. Each point corresponds to a state, with the horizontal and vertical axes showing the Shapley value for the first and second die, respectively. Points are coloured according to which dice the optimal policy re-rolls. The quadrant structure perfectly aligns with negative contributions and the need to re-roll, suggesting that value-based attributions can signal which features should be altered to improve return.
@@
Interestingly, this pattern holds across the entire state space. Figure 9 shows the Shapley values for each die in every state, with each point coloured according to which dice are re-rolled under the optimal policy. The quadrant structure reveals a perfect correspondence: a die receives a negative attribution if and only if it is re-rolled. This illustrates how value-based explanations could inform control, highlighting features that should be changed to improve return—an idea we visited in our future work discussion (Section 10.3).
B.4 Additional Example on Explaining Behaviour, Performance and Value Estimation
@@
Figure 10: Shapley value attributions for two states in Taxi (top and bottom). First column: feature contributions to the probability of selecting the optimal action taken (behaviour). Second column: contributions to the agent’s performance. Third column: contributions to the agent’s return estimate (value estimation). Fourth column: visualisation of the corresponding state. The taxi’s location is marked with a rectangle, the passenger location with a p, and the destination is circled.
Finally, we consider explaining policy, performance and value estimation in the classic Taxi domain (Dietterich, 2000). In this domain, an agent must navigate a gridworld to pick up a passenger and deliver them to their destination. Each state is described by four features: the taxi’s $(x,y)$ position, the passenger location (R, G, B, Y, or in-taxi), and the destination (R, G, B, or Y). Each step incurs a reward of $-1$, with a bonus of $+20$ for a successful drop-off and a penalty of $-10$ for inappropriate pick-up and drop-off actions. We use the OpenAI Gym Taxi-v3 implementation (Brockman et al., 2016); see there for further details. Using our proposed framework, we analyse two representative states from an optimal policy (Figure 10).
In the top state, the taxi moves south towards the passenger at location B, and the destination is location G. The behaviour explanation highlights the passenger location as the dominant contributor, with the $x$ and $y$ coordinates contributing similarly. Surprisingly, the destination receives a small negative attribution, likely because knowing it without the passenger’s location can increase the chance of taking suboptimal actions. For performance, the $x$ and $y$ coordinates contribute positively, though $x$ more than $y$, illustrating that while these features are equally important for selecting actions, their consequences for return may differ. For value estimation, the $y$ coordinate and passenger location lower the return estimate, reflecting that the agent is not in the same square as the passenger and has not picked them up.
In the bottom state, the taxi moves south to deliver the passenger, now onboard, to the destination at location B. The behaviour explanation now highlights the passenger and destination locations as similar key contributors. This illustrates the agent’s competing goals: keep the passenger in the taxi and drop them off at the destination. The performance explanation mirrors this pattern. In contrast to the top state, the passenger’s presence in the taxi—indicating a higher return—now increases the return estimate for value estimation.
This example shows how the same features influence each explanation differently depending on their role in the task and the agent’s progress toward its objective.

### v2

@@ 10.3 Using Predictions to Inform Behaviour
SVERL’s explanations of predictions reveal how feature values influence the agent’s estimated expected return. As shown in the examples in Section 7, feature values in a given state that also appear in high-return states tend to receive positive attributions. In contrast, feature values associated with lower-return states tend to receive negative attributions. This suggests that these explanations are useful for highlighting which aspects of the environment correlate with successful estimated outcomes. One intriguing possibility is to use this information to inform behaviour—for example, encouraging actions that lead to positively attributed features, or identifying negatively attributed features that should be changed to improve expected return. The additional example in Appendix B illustrates this idea concretely: in a dice-rolling domain, each die is treated as a feature and receives a negative attribution if and only if the optimal policy re-rolls the corresponding die to change its value. How to effectively incorporate these signals into decision-making or policy optimisation is a compelling question, and one we believe represents an important direction for future work.
10.4 Connecting SVERL to Broader Attribution Methods in Cooperative Game Theory
@@
SVERL provides a principled framework for explaining reinforcement learning agents using Shapley values. It introduces coalitional games tailored to reinforcement learning by defining characteristic functions for behaviour, outcomes, and predictions. Because the framework inherits the structure and axiomatic foundations of classic cooperative game theory, it can naturally accommodate alternative definitions, theoretical developments, and methodological advances from the broader literature.
To illustrate the kinds of opportunities this generality creates, we highlight two possible extensions. First, Frye et al. (2020b) propose computing Shapley values under a causal model, modifying the Shapley sum to account for known dependencies between features. For example, if one feature is a deterministic causal ancestor of another, standard Shapley values may split credit between them; a causal approach would attribute responsibility to the source feature. Extending SVERL in this way could yield explanations that more faithfully reflect the causal structure of the environment or agent behaviour.
@@
Another possibility is to explore alternative solution concepts from cooperative game theory. For instance, Banzhaf values (Banzhaf III, 1964) define a different notion of attribution based on a distinct set of axioms and may offer complementary insights when applied to SVERL’s characteristic functions. These examples demonstrate only two possible ways broader theoretical developments in attribution research can be incorporated into reinforcement learning using SVERL. Exploring these connections provides a promising avenue for future work.
@@ 11 Conclusion
Reinforcement learning agents can surpass human performance, but the how and why of their interactions with their environments remain difficult to interpret. In this work, we introduced Shapley Values for Explaining Reinforcement Learning (SVERL), a theoretical framework for explaining agent behaviour, outcomes, and predictions through the influence of features at specific states. These explanations are grounded in the well-established cooperative game-theoretic foundations of Shapley values, providing a family of theoretically principled methods for attribution.
The three explanatory elements—behaviour, outcomes, and predictions—offer distinct but complementary perspectives on agent-environment interaction. Their interdependence reveals how features can play layered roles in shaping actions, consequences, and estimates, highlighting the complexity of attribution in sequential settings and the value of unified explanatory frameworks.
SVERL formalises feature influence using Shapley values, offering principled, axiomatic explanations. However, these quantify statistical contributions rather than causal mechanisms, and thus demand careful interpretation. Translating mathematically grounded explanations into practical, trustworthy insights will require new tools, clearer user guidance, and closer alignment with findings from cognitive science.
Scaling SVERL introduces both computational and cognitive challenges. While Monte Carlo approximations can make attribution tractable, and aggregating local attributions into behavioural summaries may reduce interpretive burden, these strategies still require empirical validation and methodological refinement to support real-world deployment.
More broadly, this work situates the explanation of reinforcement learning within a rich theoretical landscape. The cooperative game-theoretic foundation invites a wide range of future extensions, including alternative characteristic functions, new solution concepts, and methodological advances, all within a common, principled language.
By establishing a theoretical basis for explaining reinforcement learning agents, SVERL contributes to the development of reinforcement learning systems that are not only powerful, but also transparent, trustworthy, and aligned with human understanding.
@@
This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) [grant number EP/X025470/1], UKRI Centre for Doctoral Training in Accountable, Responsible and Transparent AI (ART-AI) [EP/S023437/1], the EPSRC Centre for Doctoral Training in Digital Entertainment (CDE) [EP/L016540/1], and the University of Bath. This research made use of Hex, the GPU Cloud in the Department of Computer Science at the University of Bath.
@@ Appendix A Proofs of Propositions and Results
The Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ for the outcome game $({\cal F},\tilde{v}^{\pi}_{s})$ are the unique allocation of the expected return $v^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\tilde{v}^{\pi}_{s}(\mathcal{C}):2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the outcome game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\tilde{v}^{\pi}_{s}),\dots,\phi_{n}(\tilde{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
@@
The Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ for the prediction game $({\cal F},\hat{v}^{\pi}_{s})$ are the unique allocation of the predicted expected return $\hat{v}^{\pi}(s)$ across the features of state $s$ that satisfies the four axioms of fair credit assignment: efficiency, linearity, symmetry, and nullity.
Proof  For a finite set of features ${\cal F}$, the characteristic function $\hat{v}^{\pi}_{s}:2^{{\cal F}}\to\mathbb{R}$ has well-defined values for all coalitions $\mathcal{C}\subseteq{\cal F}$. Therefore, the prediction game is well-defined. By Shapley’s theorem (Shapley, 1953), the Shapley values $\phi_{1}(\hat{v}^{\pi}_{s}),\dots,\phi_{n}(\hat{v}^{\pi}_{s})$ are the unique value allocation satisfying efficiency, linearity, symmetry, and nullity.
@@
When comparing expected-return estimates using mean squared error, the function $\hat{u}^{\pi}$ is the closest in expectation to $\hat{v}^{\pi}$ among all functions defined over the subset of features $\mathcal{C}$:
@@ Appendix B Additional Examples
We present three examples to illustrate how our framework attributes behaviour, outcomes and predictions to features. We then present an additional example explaining all three.
@@
Figure 6: Explaining behaviour in a $\times 2$ 2 gridworld.
[…]
@@ 2 N 1 1 2/32/32 / 3 6/76/76 / 7
1 1/71/71 / 7 6.006.006.00 6.006.006.00 4.004.004.00 0.000.000.00 4.004.004.00 2.002.002.00
2 2/72/72 / 7 7.007.007.00 7.007.007.00 6.506.506.50 6.836.836.83 0.330.330.33 0.17-0.17- 0.17
Figure 8: Explaining expected return in a five-state gridworld.
@@
Explaining Outcomes. Consider the five-state gridworld shown in Figure 8 where the agent begins uniformly at random in state 1 or 2, receiving a reward of 1-1- 1 per action and an additional +10+10+ 10 upon reaching a goal state. Invalid actions incur a reward but leave the agent’s position unchanged. The figure shows the optimal deterministic policy $\pi$: East from state 1 and North from all others. Each state is described by two features: its $x$- and $y$-coordinates. We analyse how the expected return from each state changes as these coordinates become known, owing to the agent’s behaviour in that state shifting accordingly. These returns define SVERL’s characteristic values for explaining outcomes, while the resulting Shapley values quantify the influence of each feature. Both are shown in the bottom table of Figure 8; the top-right table shows the corresponding action probabilities under different subsets of known features.
@@
Consider state 1. When neither feature is known, the agent has probability 1/71/71 / 7 of being in state 1 (with optimal action East) and probability 2/72/72 / 7 of being in states 2, 3, and 4 (all with optimal action North). Therefore, the agent selects North with probability 6/76/76 / 7 and East with 1/71/71 / 7, yielding an expected return of 0 from state 1. If the $x$-coordinate becomes known, the agent must be in state 1 and selects East, yielding a return of 666. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability 2/32/32 / 3 and East with 1/31/31 / 3, giving an expected return of 444. This difference is reflected in the Shapley values in Figure 8, which assign twice as much influence on expected return to the $x$-coordinate as the $y$-coordinate.
@@
Now consider state 2. As in state 1, when no features are known, the agent follows the steady-state distribution, selecting North with probability 6/76/76 / 7 and East with probability 1/71/71 / 7, yielding an expected return of 6.836.836.83. If the $x$-coordinate becomes known, the agent must be in states 2, 3, or 4 (all with optimal action North). Therefore, the agent deterministically selects North, increasing the return from state 2 to 777. If instead the $y$-coordinate becomes known, the agent could be in state 1 or 2, with state 2 twice as likely, and so selects North with probability 2/32/32 / 3 and East with 1/31/31 / 3. This gives a lower expected return of 6.56.56.5 because the likelihood of the optimal action (North) has decreased from 6/76/76 / 7 to 2/32/32 / 3. The Shapley values in Figure 8 reflect this: the $x$-coordinate contributes positively to expected return, while the $y$-coordinate has a negative influence.
@@
(3,6)(3,6)( 3 , 6 ) 0.0240.0240.024 0.670.670.67 0.450.450.45 0.900.900.90 0.660.660.66 0.22-0.22- 0.22 0.230.230.23
(1,1)(1,1)( 1 , 1 ) 0.0180.0180.018 0.360.360.36 0.450.450.45 0.450.450.45 0.660.660.66 0.15-0.15- 0.15 0.15-0.15- 0.15
Table 7: Explaining predictions in a dice-rolling game.
@@
Explaining Predictions. Consider a dice-rolling game, in which each episode begins with a random roll of two six-sided dice. State is denoted $(d_{1},d_{2})$, where $d_{1}$ and $d_{2}$ are the first and second die values, respectively. The agent may re-roll either die, both dice, or neither. After each action, the episode ends with probability 0.50.50.5, yielding a reward of 111 if the sum of the dice is 101010 or more, and 0 otherwise. The optimal policy keeps 555s and 666s, and re-rolls 111s, 222s, 333s, and 444s. We analyse how the agent’s predicted expected return in each state changes as one or both dice become known. These returns define the characteristic values in SVERL’s prediction game. The resulting Shapley values quantify each die’s contribution, as shown in Table 7.
@@
Consider state (3,6)(3,6)( 3 , 6 ). When neither die is known, the value estimate is the average return over the steady-state distribution: 0.660.660.66. If $d_{2}$ becomes known to be 666, the estimate increases to 0.900.900.90 because states with a 666 have higher average returns. If instead $d_{1}$ becomes known to be 333, the estimate drops to 0.450.450.45 because states with a 333 have lower average returns. The Shapley values reflect this difference: the 666 contributes positively to the value estimate, while the 333 contributes negatively. Now consider state (1,1)(1,1)( 1 , 1 ). As above, when neither die is known, the estimate is 0.660.660.66. If either die becomes known to be 111, the estimate decreases to 0.450.450.45 because states with a 111 have a lower average return. When both dice are known, the agent must be in the lowest-return state, and the estimate falls further to 0.360.360.36. Thus, each die lowers the value estimate by the same amount, as reflected in their equal negative Shapley values. Interestingly, this pattern holds across the entire state space. Figure 9 shows the Shapley values for each die in every state, with each point coloured and shaped according to which dice are re-rolled under the optimal policy. The quadrant structure reveals a perfect correspondence: a die receives a negative attribution if and only if it is re-rolled. This illustrates how prediction-based explanations could inform control, highlighting features that should be changed to improve predicted return.
Figure 9: Shapley values for the value estimates of every state in the dice-rolling game. Each point corresponds to a state, with the horizontal and vertical axes showing the Shapley value for the first and second die, respectively. Points are coloured according to which dice the optimal policy re-rolls.
@@
Figure 10: Shapley values for two states in Taxi (top and bottom), showing the feature contributions to the probability of selecting the optimal action taken (first column), to the agent’s expected return (second column), and the agent’s estimated expected return (third column). On the right-hand side, the corresponding taxi state is shown, with the taxi location marked by a rectangle, the passenger location with a p, and the destination with a circle.
Explaining Behaviour, Outcomes and Predictions. We use the classic Taxi domain (Dietterich, 2000) to show how the same features influence each explanation differently depending on their role in the task and the agent’s progress toward its objective. In this domain, an agent navigates a gridworld, picking up a passenger and delivering them to their destination. Each state is described by four features: the taxi’s $(x,y)$ position, the passenger location (R, G, B, Y, or in-taxi), and the destination (R, G, B, or Y). Each step incurs a reward of 1-1- 1, with a bonus of +20+20+ 20 for a successful drop-off and a penalty of 10-10- 10 for inappropriate pick-up and drop-off actions. We use the OpenAI Gym Taxi-v3 implementation (Brockman et al., 2016). We analyse an optimal policy at two states (Figure 10).
In the top state, the taxi moves south towards the passenger at location B, and the destination is location G. The behaviour explanation highlights the passenger location as the dominant contributor, with the $x$ and $y$ coordinates contributing similarly. Surprisingly, the destination receives a small negative attribution, possibly because knowing it without knowing the passenger’s location can increase the probability of taking suboptimal actions towards it. For outcome, the $x$ and $y$ coordinates contribute positively, though $x$ more than $y$, illustrating that while these features are similarly important for selecting actions, their consequences for return may differ. For prediction, the $y$ coordinate and passenger location lower the predicted expected return, reflecting that the agent is not in the same row as the passenger and has not picked them up.
In the bottom state, the taxi moves south to deliver the passenger, now onboard, to the destination at location B. The behaviour explanation now highlights the passenger and destination locations as similar key contributors. This illustrates the agent’s competing goals: keep the passenger in the taxi and drop them off at the destination. The outcome explanation mirrors this pattern. In contrast to the top state, the passenger’s presence in the taxi—indicating a higher return—now increases the predicted expected return.